OpenStudy (amistre64):

inital conditions? \(f'(x)=\cfrac{x^2-1}{x}\): f(1)= 1/2, f(-1) = 0

OpenStudy (angela210793):

u have to integrate this right?

OpenStudy (amistre64):

yep

OpenStudy (angela210793):

i mean in order to get f(x)

OpenStudy (anonymous):

I can do the integration Job \[f'(x) = x - \frac{1}{x}\] \[f(x)=\frac{x^2}{2} + log_ex + C\]

OpenStudy (amistre64):

good, now determine C by the inital conditions :)

OpenStudy (anonymous):

\[f(x)=\frac{x^2}{2} - log_ex + C\]

OpenStudy (angela210793):

f(x)=x^2/2-lnx+c then by f(1) and f(-1) u find c

OpenStudy (anonymous):

\[f(1) = \frac{1}{2} + C = \frac{1}{2}\]

OpenStudy (anonymous):

\[C = 0\]

OpenStudy (amistre64):

i can get it for f(1) just fine ..., now try it with f(-1) = 0

OpenStudy (amistre64):

when does ln|x| + C = 1/2 AND 0?

OpenStudy (anonymous):

Okay, Lol I see it now.

OpenStudy (amistre64):

:)

OpenStudy (amistre64):

this is a homework problem straight from the textbook ... and i gotta wonder

OpenStudy (angela210793):

Gosh...i hate everything tht includes integrals :/

OpenStudy (amistre64):

C can be anything so I even tried: ln(x) + ln(C) = ln(Cx) ln(x) - ln(C) = ln(x/C)

OpenStudy (zarkon):

why are there two initial conditions? are you sure you wrote the problem correctly?

OpenStudy (zarkon):

oh yea...

OpenStudy (amistre64):

im sure .... im wondering if it aint spose to be a double derived

OpenStudy (zarkon):

so you have a +c1 for x<0 and a +c2 for x>0

OpenStudy (zarkon):

piecewise function

OpenStudy (amistre64):

.... i gotta piece it in? thats soo feels like cheating lol

OpenStudy (anonymous):

two different constants!... First time for me...

OpenStudy (zarkon):

\[f(x)=\left\{\begin{matrix}\frac{x^2}{2} -\ln|x| + C_1 & \text{ if } x>0 \\ \frac{x^2}{2} - \ln|x| + C_2 & \text{ if } x<0 \end{matrix}\right.\]

OpenStudy (amistre64):

yep, the answer key for the one like it shows a piecewise as well ....

OpenStudy (amistre64):

none of the examples in the chapter even give a hint to that ....

OpenStudy (zarkon):

this "it is +c on evey open interval that is continiuous" is probably buried somewhere in the text

OpenStudy (zarkon):

most of the time we ignore this

OpenStudy (amistre64):

yeah, most of the time :) thnx for the help yall

OpenStudy (zarkon):

sorry...not continuous...i mean defined

OpenStudy (zarkon):

your function is defined on (-infinity,0) and (0,infinity) so we will have two constants

OpenStudy (amistre64):

\[f(x)=\left\{\begin{matrix}\frac{x^2}{2} -\ln|x| & \text{ if } x>0 \\ \frac{x^2}{2} - \ln|x| -\frac{1}{2} & \text{ if } x<0 \end{matrix}\right.\]looks good then